sb CH 8 A company purchased three identical inventory items for $800, $900, and $1,120 in...

sb CH 8

A company purchased three identical inventory items for $800, $900, and $1,120 in that order. It then sold 1 of the units for $1,500. Assuming the company uses the average-cost method for inventory the gross profit
Answer #1

Under the average cost method, first need to compute the average cost per unit. Then multiply the average cost per unit to ending units to get ending inventory and multiply with units sold to get cost of goods sold.

In the above question,

Total units available for sale = 3 units.

Total cost of units available for sale = 800 + 900 + 1,120

Total cost of units available for sale = $ 2,820.

Average cost per unit = 2,820/3

Average cost per unit = $ 940 per unit.

Cost of goods sold = 940 x 1

Cost of goods sold = 940.

To find out the gross profit,

Gross profit = Sales - Cost of goods sold.

Gross profit = 1,500 - 940

Gross profit = $ 560.

Hence, $ 560 is the correct answer.

SUMMARY:

Multiplying average cost per unit.with units sold resulting cost of goods sold. Sales value is given. So difference if these will give gross profit.

Hence, $ 560 is the correct answer.